OpenStudy (solomonzelman):

I'm looking at the following series (for fun), and I this is what I got....

OpenStudy (solomonzelman):

\(\large\color{#000000}{ \displaystyle \sum_{ n=2 }^{ \infty } ~ \frac{ 1}{\ln\Gamma(kx)}}\) When \(k<0\) and \(k\notin \mathbb{Z}\) I did the limit test and the limit (and thus the series) diverges (does not approach a value). (I know if it converged, no colnclusions...) When \(k\le0\) and \(k\in \mathbb{Z}\) Then all terms are zeros and series converges.

OpenStudy (solomonzelman):

\(\color{#000000}{ \displaystyle \sum_{ n=2 }^{ \infty } ~ \frac{ 1}{\ln\Gamma(kx)}}\) When \(k\in \mathbb{N}\). We can show divergence by the comparison Test. \(\color{#000000 }{ \displaystyle \frac{1}{\ln\Gamma (kx)}>\frac{1}{\ln (kx{\tiny~}!)}>\frac{1}{\ln \left[(kx)^{kx}\right] } }\) \(\color{#000000 }{ \displaystyle \frac{1}{\ln\Gamma (kx)}>\frac{1}{kx\ln \left[kx\right] } }\) Therefore, \(\color{#000000}{ \displaystyle \sum_{ n=2 }^{ \infty } ~ \frac{ 1}{\ln\Gamma(kx)}>\sum_{ n=2 }^{ \infty } ~ \frac{1}{kx\ln \left[kx\right] }}\) And since the series on the right diverges. (When integrated, I got (1/k)•ln\((\)ln(kx)\()\), and the limit of this as \(x\to\infty\) will diverge to ∞.)

OpenStudy (solomonzelman):

When \(k\notin \mathbb{N}\) and \(k>0\). We can show that the series converges by the squeeze theorem. For example, we know that \(\color{#000000}{ \displaystyle \sum_{ n=2 }^{ \infty } ~ \frac{ 1}{\ln\Gamma(\color{red}{100}x)}}\) and \(\color{#000000}{ \displaystyle \sum_{ n=2 }^{ \infty } ~ \frac{ 1}{\ln\Gamma(\color{red}{101}x)}}\) both diverge. Therefore, (for instance), \(\color{#000000}{ \displaystyle \sum_{ n=2 }^{ \infty } ~ \frac{ 1}{\ln\Gamma(\color{red}{100}x)}>\sum_{ n=2 }^{ \infty } ~ \frac{ 1}{\ln\Gamma(\color{blue}{100.6}x)}>\sum_{ n=2 }^{ \infty } ~ \frac{ 1}{\ln\Gamma(\color{red}{101}x)}}\) And since the left and right diverge (because they are just the same case of \(k\in\mathbb{N}\)), therefore the middle series diverges by the Squeeze Theorem.

OpenStudy (solomonzelman):

For \(0<k<1\), and if \(k\notin\mathbb{Z}\). We can use the Comprison Test as well. \(\color{#000000 }{ \displaystyle \frac{1}{\ln\Gamma (kx)}>\frac{1}{\ln \left[(kx)^{kx}\right] } }\) (This would be true if not at n=2, then for sure from some n=j and on, and if we show that series diverges from n=j then it would diverge from n=2 too) So, in this case; \(0<k<1\), and if \(k\notin\mathbb{Z}\), series diverges.

OpenStudy (benlindquist):

wt

OpenStudy (anonymous):

umm....

OpenStudy (anonymous):

hes either a rocket scientist or a whale bioligist(i know what im doing im a whale bioligist)

OpenStudy (anonymous):

Maybe a silly question, but since your index is n, I'm assuming you intentionally do not have an n in the summand, or is it missing? If we are not supposed to have n in the summand, then we're really only considering \[\lim_{n \rightarrow \infty} \frac{ n }{ \ln \Gamma (kx) }\] Am I missing something?

OpenStudy (solomonzelman):

Yes, I apologize, that is supposed to be nx.

OpenStudy (solomonzelman):

Or else the unknown constant is the product of k and x, rather than just k. Thanks for asking.

OpenStudy (anonymous):

yep whale bioligist